OpenStudy (anonymous):

(Thermochemistry) A 200 gallon water bed is filled with water that is initially at 18 degrees C. How many kcal will it take to heat the bed to 70 degrees C?

OpenStudy (anonymous):

First find mass of the water present in 1 gallon assuming 1kg/m^3 density. Then apply the equation Q=MC deltaT Here C for water =4200J/kgK

OpenStudy (anonymous):

I did not understand any of what you just said other than that I should find the mass of water present in one gallon.

OpenStudy (anonymous):

Ok so you understood how to find mass?

OpenStudy (anonymous):

No. Only that I should find it.

OpenStudy (anonymous):

Ok density of water is 1kg per metre cube ok?

OpenStudy (anonymous):

Ok. Is that is just the units we're using here?

OpenStudy (anonymous):

denisty=mass/volume. mass= density*volume. If i'm not mistaken, 1 gallon=3.79 Litres or 0.00379 Metre cube density =1 so mass=0.00379kg or 3.79 g

OpenStudy (anonymous):

Now we know mass of water that is there, Do you know about a quantity called specific heat capacity?

OpenStudy (anonymous):

oops wait there are 200 gallons so multiply by 200. mass=3.79*200=758g

OpenStudy (anonymous):

sorry for the confusion there are 200 gallons and 1gallon=0.00379 m^3 so mass=1*200*0.00379 or 0.758kg.or 758g. ok?

OpenStudy (anonymous):

Ok, I've got you.

OpenStudy (anonymous):

ok so do you know specific heat capacity?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok Good you define it as amount of heat required to raise 1kg of anything by 1degree celsius ok?

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

ok so it follows that to find heat to raise 0.758 kg by 52 degree celsius you have to just multiply specific heat*52*0.758. Specific heat of water in kilo calories is just 1. so heat=52*0.758=39.416 Kcal

OpenStudy (anonymous):

Did you get it?

OpenStudy (anonymous):

Yes, I think I do! Thanks!

OpenStudy (anonymous):

You're welcome :)